# Gaussian Head Avatar: Ultra High-fidelity Head Avatar via Dynamic Gaussians

### v2

@@ 5 Experiments
Figure 3: Qualitative comparisons of different methods on self reenactment task. From left to right: NeRFBlendShape [17], NeRFace [16], HAvatar [71] and Ours. Our method can reconstruct details like beards, teeth, etc. with high quality.
@@ 5.1 Implementation Details
In the experiment we use 12 sets of data, 10 of which are from NeRSemble [28], and the other 2 are multi-view video data from HAvatar [71]. For the 10 identities from NeRSemble, each set contains 2500 to 3000 frames, 16 cameras are distributed about 120 degrees in front, and simultaneously capture 2K resolution video. For each identity, We use the sequences marked with ”FREE” as the evaluation data, and the rest as the training data. For the 2 identities from HAvatar, each set contains 3000 frames, 8 cameras are distributed about 120 degrees in front, and 4K resolution videos are collected simultaneously. Later, we crop the face area and resize to 2K resolution.
For data preprocessing, we first remove the background [35] and extract 68 2D facial landmarks [10] for all the images. Then, for each frame, we use multi-view images to estimate the corresponding 3D landmarks, expression coefficients and head pose by fitting the Basel Face Model (BFM) [18] to the extracted 2D landmarks. Note that we define the 3D landmarks as the usual 68 landmarks with vertices indexed as multiples of 100 in the BFM vertices.
@@
Table 1: Quantitative evaluation results of NeRFBlendShape [17], NeRFace [16], HAvatar [71], our method without super resolution and our full method on self reenactment task.
@@
Table 2: Quantitative evaluation results of the other SOTA methods and our method on 3D consistency.
Figure 5: Novel view synthesis results of our method. We use 8-view synchronized videos for training the avatar.
@@ 5.2 Results and Comparisons
Next, we conduct a quantitative evaluation for the four methods on 5 identities and 6 cameras using the evaluation split. The evaluation metrics include: Peak Signal-to-Noise Ratio (PSNR), Structure Similarity Index (SSIM), Learned Perceptual Image Patch Similarity (LPIPS) [70] and Fréchet Inception Distance (FID) [22]. Note, we calculate FID by comparing the distribution of all the training images and all the rendered images. As the task mainly focuses on the reconstruction of the head, we use face parsing  [^1] https://github.com/zllrunning/face-parsing.PyTorch to remove the body parts in the image to eliminate their impact in the experiment. As shown in Tab. 1, our method demonstrates a slight improvement in PSNR and SSIM compared with previous methods, and a significant improvement in LPIPS and FID, which means that our method can generate more high-frequency details.
@@
Cross-Identity Reenactment. We also qualitatively compare our method with the above SOTA methods on cross-identity reenactment task. As shown in the results, our method is able to synthesize higher-fidelity images with more accurate expression transfer and richer emotions.
@@
Novel View Synthesis. In this section, we show the results of novel view synthesis as shown in Fig. 5 shown. In this case, we use video data from 8 views for training and render the image at a new viewpoint. Next, we quantitatively evaluate the 3D consistency of our method and compare it with other SOTA methods mentioned above. Specifically, we select the 5 identities as above and use the video data from 8 cameras for training while the other 8 hold-out cameras for evaluation. The evaluation metrics include PSNR, SSIM and LPIPS (512 resolution). The results are shown in Tab. 2. Our method outperforms other methods in 3D consistency.
@@ 5.3 Ablation Study
Table 3: Quantitative evaluation results of the other two ablation baselines and ours on self reenactment task.
Ablation on Initialization Strategies. In order to verify the effectiveness of our geometry-guided initialization strategy 4.3, we compare it with the strategy to use the FLAME model for initialization (FLAME-Init). Specifically, after fitting a FLAME model through multi-view data, we first subdivide the FLAME mesh 4 times and use the neutral vertices as the positions of the neutral Gaussians. Then, the expression deformation MLP is optimized to learn the displacement of FLAME vertices. We set the per-vertex feature to zeros, while randomly initialize the parameters of the expression color MLP. The initialization of other variables is the same as our strategy. Qualitative results are shown in the Fig. 6. Due to the lack of initialization for the hair and shoulders in FLAME-Init, the points to model these parts are offset from nearby vertices, which leads to sparseness of the Gaussians, resulting in blurring.
@@
Ablation on Deformation Modeling Approaches. We compare our fully learned deformation field with the previous mesh-based deformation (Mesh-Deform). Specifically, we migrate the method in INSTA [76] for controlling the NeRF deformation to our Gaussians. First we fit a 3DMM mesh template. Then, for each Gaussian point, find the closest face on the mesh, and calculate the deformation gradient to estimate the displacement. Qualitative results are shown in the Fig. 7. For some expressions that cannot be captured well by the 3DMM mesh template, our method can learn accurate deformation, thereby achieving the modeling of complex expressions. Quantitave results are shown in the Fig. 3. Our method outperforms both the two ablation baselines on PSNR, SSIM and LPIPS metrics.
@@ 6 Discussion and Conclusion
Limitation. For the tongue and teeth inside the mouth or long hair, blurring is sometimes produced in our method due to the lack of tracking methods.
Conclusion. In this paper, we propose Gaussian Head Avatar, a novel representation for head avatar reconstruction, which leverages dynamic 3D Gaussians controlled by a fully learned expression deformation. Experiments demonstrate our method can synthesize ultra high-fidelity images while modeling exaggerated expressions. In addition, we propose a well-designed minute-level initialization strategy to ensure the training convergence. We believe our Gaussian Head Avatar will become the mainstream direction for head avatar reconstruction in the future.
Acknowledgement. This paper is supported by National Key R&D Program of China (2022YFF0902200), the NSFC project No.62125107, the Beijing Municipal Science & Technology Z231100005923030.

### v3

@@ 4 Experiments
Figure 4: Qualitative comparisons of different methods on self reenactment task in NeRSemble dataset [5]. From left to right: NeRFBlendShape [70], NeRFace [68], HAvatar [5] and Ours. Our method can reconstruct details like beards, teeth, eyes, etc. with high quality.
Figure 5: Qualitative comparisons of different methods on self reenactment task with dynamic hairs in the self-captured dataset. From left to right: HAvatar [5], GaussianAvatars [13], MeGA [98] and Ours. Our method can reconstruct details with high quality.
@@ 4.1 Implementation Details
In the experiment, we use 15 sets of data, with 10 from NeRSemble [25], 2 from multi-view video data from HAvatar [5], and 3 extra self-captured and constructed data containing hair motion as there is no complete non-rigid dynamic hair trajectory (e.g., nodding and swinging at different speeds) in NeRSemble [25]. For the 10 identities from NeRSemble, each set contains 2500 to 3000 frames, 16 cameras are distributed about 120 degrees in front, and simultaneously capture 2K resolution video. For each identity, We use the sequences marked with ”FREE” as the evaluation data, and the rest as the training data. For the 2 identities from HAvatar, each set contains 3000 frames, 8 cameras are distributed about 120 degrees in front, and 4K resolution videos are collected simultaneously. Later, we crop the face area and resize all images to 2K resolution. For the 3 identities from self-captured data, each set contains about 1000 frames, 4 cameras are distributed about 90 degrees in front, and 4K resolution videos are collected simultaneously. Later, we crop the face area and resize all images to 2K resolution.
For data preprocessing, we first remove the background [94], segment the hair from the head by the face-parsing [95], and extract 68 2D facial landmarks [99] for all the images. Then, for each frame, we use multi-view images to estimate the corresponding 3D landmarks, the expression coefficients, and the head pose by fitting the Basel Face Model (BFM) [65] to the extracted 2D landmarks. Note that we define the 3D landmarks as the usual 68 landmarks with vertices indexed as multiples of 100 in the BFM vertices.
4.2 Training Details
During the geometry-guided initialization stage, we use an Adam optimizer, and set the learning rate to $1\times 10^{-3}$ for all the networks and $1\times 10^{-4}$ for the neutral 3D landmarks $\boldsymbol{P}_{0}$. Then, we train the model for 10000 iterations with a batch size of 4. During the Gaussian model training stage, we also use an Adam optimizer, and set the learning rate to $1\times 10^{-4}$ for the three color MLPs $\{\boldsymbol{f}_{col}^{exp},\boldsymbol{f}_{col}^{pose},\boldsymbol{f}_{col}^{hair}\}$, the three deformation MLPs $\{\boldsymbol{f}_{def}^{exp},\boldsymbol{f}_{def}^{pose},\boldsymbol{f}_{def}^{hair}\}$, and the three attribute MLPs $\{\boldsymbol{f}_{att}^{exp},\boldsymbol{f}_{att}^{pose},\boldsymbol{f}_{att}^{hair}\}$, $1\times 10^{-5}$ for the neutral positions $\boldsymbol{X}_{0}$ and the point-wise feature vectors $\boldsymbol{F}_{0}$, $1\times 10^{-4}$ for the neutral rotation $\boldsymbol{Q}_{0}$, $3\times 10^{-4}$ for the neutral scale $\boldsymbol{S}_{0}$, $1\times 10^{-3}$ for the neutral opacity $\boldsymbol{Q}_{0}$ and $1\times 10^{-4}$ for the super resolution network $\boldsymbol{\Psi}$. Finally, we train the Gaussian model for 600000 iterations with a batch size of 1 until fully convergence.
@@ 4.2 Training Details
TABLE I: Quantitative evaluation results of NeRFBlendShape [70], NeRFace [68], HAvatar [5], our method without super resolution and our full method on self reenactment task in NeRSemble dataset [5] and HAvatar dataset [5].
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
HAvatar 26.25 0.879 0.063 36.34
GaussianAvatars 24.21 0.823 0.181 65.27
MeGA 25.13 0.880 0.165 55.34
Ours 27.05 0.883 0.060 30.53
TABLE II: Quantitative evaluation results of the other SOTA methods and our method on self reenactment task with the motion of the hair in our self-captured video.
@@
Figure 7: Qualitative comparisons of different methods on cross-identity reenactment task with dynamic hairs in the self-captured dataset. From left to right: HAvatar [5], GaussianAvatars [13], MeGA [98] and Ours. Our method synthesizes high-fidelity images while ensuring the accuracy of hair movement.
@@
TABLE III: Quantitative evaluation results of the other SOTA methods and our method on 3D consistency.
Figure 8: Novel view synthesis results of our method. Top: we use 8-view synchronized videos for training the avatar. Bottom: we use 4-view synchronized videos for training the avatar with dynamic hais.
@@ 4.3 Results and Comparisons
Next, we conduct a quantitative evaluation for the four methods on 5 identities and 6 cameras using the evaluation split. The evaluation metrics include: Peak Signal-to-Noise Ratio (PSNR), Structure Similarity Index (SSIM), Learned Perceptual Image Patch Similarity (LPIPS) [96] and Fréchet Inception Distance (FID) [100]. Note, we calculate FID by comparing the distribution of all the training images and all the rendered images. As the task mainly focuses on the reconstruction of the head, we use face-parsing [95] to remove the body parts in the image to eliminate their impact in the experiment. As shown in Tab. I, our method demonstrates a slight improvement in PSNR and SSIM compared with previous methods, and a significant improvement in LPIPS and FID, which means that our method can generate more high-frequency details. Note that HHAvatar only made additional changes to the position and attributes of the Gaussians in the hair part, and the final rendering method remains the same as the previous version[33]. Therefore, when the dataset is organized chronologically and the hair maintains relatively static compared to the head, the reconstructed results are consistent with the previous version[33]. When the dataset is not in chronological order, simply canceling the change in hair inertia will result in the same reconstruction as the previous version[33].
We also compare our method with existing SOTA methods in qualitative experiments on self reenactment task with dynamic hairs. Qualitative results on self reenactment task are shown in the Fig. 5. It can be seen that all other methods fail to capture the accurate physical motion of the hair component and the hair part is blurry. Furthermore, as a result of the occlusion created by hair during motion, the face itself appears blurry. Because all existing methods neglect the kinematic attributes of hair, none of them can successfully complete self reenactment task with the motion of the hair. By modeling hair separately from the human head and introducing the kinematic features of the hair, our method can accurately reconstruct high-frequency details of the hair. Besides, our method can capture the realistic physical motion of hair.
Quantitative results on self reenactment task with dynamic hairs are shown in Tab. II. As shown in Tab. II, our method demonstrates a slight improvement in PSNR and SSIM compared with previous methods, and a significant improvement in LPIPS, which means that our method can generate more high-frequency details.
@@
Cross-Identity Reenactment. We qualitatively compare our method with the above SOTA methods on cross-identity reenactment task. As shown in Fig. 6, our method is able to synthesize higher-fidelity images with more accurate expression transfer and richer emotions.
We also qualitatively compare our method with the above SOTA methods on cross-identity reenactment task with dynamic hairs. As shown in Fig. 7, other methods did not consider the status of the hair at the previous moment, resulting in no reasonable physical motion of the hair during driving. Therefore, those methods fail to capture the accurate physical motion of the hair component and the hair part is blurry. Our method is able to synthesize higher-fidelity images with more accurate hair motion.
@@
Novel View Synthesis. In this section, we show the results of novel view synthesis as shown in the top of Fig. 8 shown. In this case, we use video data from 8 views for training and render the image at a new viewpoint. Next, we quantitatively evaluate the 3D consistency of our method and compare it with other SOTA methods mentioned above. Specifically, we select the 5 identities as above and use the video data from 8 cameras for training while the other 8 hold-out cameras for evaluation. The evaluation metrics include PSNR, SSIM and LPIPS (512 resolution). The results are shown in Tab. III. Our method outperforms other methods in 3D consistency.
We also show the results of novel view synthesis with dynamic hairs, and the results are shown in the bottom of Fig. 8. In this case, we use our self-captured video data from 4 views for training and render the image at five new viewpoints. As shown in the bottom of Fig. 8, our method can maintain 3D consistency and synthesize high-fidelity images with accurate hair motion.
@@ 4.4 Ablation Study
TABLE IV: Quantitative evaluation results of the other two ablation baselines and ours on self reenactment task.
Ablation on Initialization Strategies. In order to verify the effectiveness of our geometry-guided initialization strategy 3.3, we compare it with the strategy to use the FLAME model for initialization (FLAME-Init). Specifically, after fitting a FLAME model through multi-view data, we first subdivide the FLAME mesh 4 times and use the neutral vertices as the positions of the neutral Gaussians. Then, the expression deformation MLP is optimized to learn the displacement of FLAME vertices. We set the per-vertex feature to zeros, while randomly initialize the parameters of the expression color MLP. The initialization of other variables is the same as our strategy. Qualitative results are shown in the Fig. 9. Due to the lack of initialization for the hair and shoulders in FLAME-Init, the points to model these parts are offset from nearby vertices, which leads to sparseness of the Gaussians, resulting in blurring.
@@
Ablation on Deformation Modeling Approaches. We compare our fully learned deformation field with the previous mesh-based deformation (Mesh-Deform). Specifically, we migrate the method in INSTA [23] for controlling the NeRF deformation to our Gaussians. First we fit a 3DMM mesh template. Then, for each Gaussian point, find the closest face on the mesh, and calculate the deformation gradient to estimate the displacement. Qualitative results are shown in the Fig. 10. For some expressions that cannot be captured well by the 3DMM mesh template, our method can learn accurate deformation, thereby achieving the modeling of complex expressions. Quantitative results are shown in the Fig. IV. Our method outperforms both the two ablation baselines on PSNR, SSIM and LPIPS metrics.
Figure 11: Ablation study on the dynamic hair module: our approach can capture the motion of the hair and reconstruct more accurate hair.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
W/o Dynamic Hair 26.23 0.825 0.123
W/o Occlusion Perception 26.83 0.874 0.116
Full Model 27.02 0.876 0.108
TABLE V: Quantitative evaluation results of the baseline and our full model on self reenactment task.
Figure 12: Ablation study on the dynamic hair module: our approach can capture the motion of the hair and reconstruct more accurate hair.
Ablation on Hair Dynamic Module. To assess the effectiveness of independently modeling hair, we designed an experiment in which we kept modeling the head and the hair in the same manner. Qualitative results are shown in the Fig. 11. It can be seen that if the hair is not modeled separately from the head, not only will the hair not have realistic physical motion, but it will also produce blurring. Therefore, it is necessary to model hair separately. Quantitative results are shown in the Tab. V. It can be seen that our method outperforms both the two ablation baselines on PSNR, SSIM and LPIPS metrics.
Ablation on Occlusion Perception Module. To assess the effectiveness of the mask that considers occlusion, we designed an experiment in which we used alpha rendering  [^2] https://github.com/ashawkey/diff-gaussian-rasterization for separating the hair and the head. Qualitative results are shown in the Fig. 12. It can be seen that due to the lack of consideration for occlusion, the hole may sometimes occur in the face during driving because of the hair motion. Therefore, it is necessary to consider the occlusion. Quantitative results are shown in the Tab. V. It can be seen that our method outperforms both the two ablation baselines on PSNR, SSIM and LPIPS metrics.
Figure 13: Failure case: our method produce relatively less exaggerated results.
@@ 5 Discussion and Conclusion
Limitation. For the tongue and teeth inside the mouth, blurring is sometimes produced in our method due to the lack of tracking methods. On the other hand, the reconstructed head avatar cannot make expressions other than those in the training set. Therefore, when the actor’s expression is too exaggerated, our method will output relatively less exaggerated results as shown in Fig. 13. When the head movement speed in the training set is excessively high, leading to complete hair coverage over facial features, the fitting of 3DMM expression coefficients will fail. Hence, it is essential to avoid moving the head too rapidly.
Conclusion. In this paper, we propose HHAvatar, a novel representation for head avatar reconstruction, which leverages dynamic 3D Gaussians controlled by a fully learned expression deformation and models the physical motion of the hairs. Experiments demonstrate our method can synthesize ultra high-fidelity images while modeling exaggerated expressions and capturing the realistic physical motion of the hair. In addition, we propose a well-designed minute-level initialization strategy to ensure the training convergence and a temporal module to fuse time series information. We believe our HHAvatar will become the mainstream direction for head avatar reconstruction in the future.